OpenStudy (anonymous):

What is the slope of the line passing through the points (2,7) and (-1,3) ? 2/7 3/4 4/3 1/3

OpenStudy (alexandervonhumboldt2):

slope formula=(y2-y1)/(x2-x1) or (y1-y2)/(x1-x2) where points are (x1, y1) and (x2, y2)

OpenStudy (anonymous):

The answer is 4/3. Slope is m= (y2-y1)/(x2-x1). 3 is y2 and 7 is y1; -1 is x2 and 2 is x1. The answer is -4/-3 which is the same as 4/3.

OpenStudy (anonymous):

omg thank you so much, I was so lost. very behind in math rn